Circular No. 7915

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002df IN MCG –01-53-6 M. Papenkova and W. D. Li, University of California at Berkeley, re-

port the LOTOSS (cf. IAUC 7906) discovery of an apparent supernova on unfiltered CCD images taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on June 2.5 (mag ∼ 17.4) and 3.5 UT (mag ∼ 17.1). SN 2002df is located at α = 20^{h}45^{m}39^{s}.33, δ = −4^{o}57′04′′.7 (equinox 2000.0), which is 27′′.1 west and 4′′.2 south of the nucleus of MCG –01-53-6. A KAIT image of the same field on May 10.5 showed nothing at this position (limiting mag ∼ 18.5).

SUPERNOVA 2002dg Further to IAUC 7905, W. M. Wood-Vasey, G. Aldering, and P. Nugent

report the discovery of another apparent supernova (mag 18.5) on unfiltered NEAT images taken with the Palomar 1.2-m Schmidt telescope on May 31.3 and June 3.2 UT. SN 2002dg is located at α = 14^{h}57^{m}16^{s}.29, δ = +5^{o}54′00′′.6 (equinox 2000.0), which is 4′′ east of the apparent host galaxy. No object is apparent at this position on NEAT images from 2002 May 5 (limiting mag 21.5; S/N = 3) or 2001 July 21 (limiting mag 20.5).

SUPERNOVA 2002db IN NGC 5683 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002db (cf. IAUC 7906), obtained by M. Calkins on June 1.29 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova near maximum. Adopting the NASA/IPAC Extragalactic Database (NED) recession velocity of 10859 km/s for the host galaxy, the supernova expansion velocity is ∼ 12200 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova was 2 ± 2 days before maximum light.

SUPERNOVA 2002de IN NGC 6104 Matheson et al. also report that a spectrum (range 370–750 nm) of SN

2002de (cf. IAUC 7914), obtained by Calkins as above on June 2.23 UT, shows it to be a type-Ia supernova before maximum. Adopting the NED recession velocity of 8428 km/s for the host galaxy, the supernova expansion velocity is ∼ 12000 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova was 4 ± 2 days before maximum light.

2002 June 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green